The OS/2 port of bind 4.9.3

Peter Meerwald

1. Juli 1996

1 What is it?

This is release 4.9.3 of bind^{1} ported to OS/2. It relies on EMX^{2} 0.9b fix 5 (or newer).

2 Installation

To install just take a look a the standard bind documentation^{3}. The directory structure for the OS/2 port is laid down in named/pathnames.h. A re-compile is neccassary to customize directory names – sorry. The EMX runtime and IBM TCP/IP have to be installed.

You will find the executables in the source directories (named/, tools/, tools/nslookup), make install does (of course) not work (yet).

It is a good idea to have syslogd running the first time you start named. Try parameter -d for extensive debug info and/or -q for query logging.

3 OS/2 specific stuff

3.1 What happened to the code?

Actual code is pretty unchanged, I adapted the makefiles, figured out a suitable compat/lib/lib44bsd.a, read conf/protability.h very carefully :) and while doing so my OS/2 box turned into a unix-lookalike^{4}.

Look for EMX defines in the code (you won’t find many) to see what changed.

3.2 Re-compiling

Especially named/Makefile contains one bad hardcoded hack (I was unable get sed to work), I used tcsh 6.0.6.1 and bash 1.12, to me, it seems that these shells lack some features :) - well I’m very lost when it comes to shell setup/shell features - I am awaiting your hints…

News flash: Try ksh 5.27 to re-compile – works much better that tcsh and bash! Compiles the entier project in one pass, wow…

In order to re-compile the bind port you’ll need a syslog library (on hobbes^{5} for IBM’s CSet, a port of the port to EMX is available from the author) and flex.

I removed the contrib/, shres/, doc/rfc/ directory from the original distribution, some postscript files about DNS security were also removed.

1Berkley Internet Name Domain

2Eberhard Mattes eXtentions – a distribution of GCC for OS/2

3look at http://www.vix.com/isc/bind.html for a reference manual (BOG – Bind Operations Guide), FAQ, protocol definitions, …

4you may wish to get GNU shell—file—text utilities from hobbes

5point your browser to ftp://ftp-os2.nmsu.edu/os2 or ftp://ftp-os2.cdrom.com

3.3 Sending signals

Unix background daemons receive commands via signals that are sent to their pid^{6}. bind 4.9.3 understands a number of signals: SIGINT, SIGQUIT, SIGIOT, SIGUSR1, SIGUSR2, SIGHUP, SIGALARM, SIGTERM.

A way to send signals under OS/2 is provided in the Apache httpd port^{7} in directory emphApache/os2/.

4 History of the OS/2 port

• June 14, 1996 — released bind493a.zip

• July 3,1996 — bind493b.zip contains this wonderful L^{A}TEX -README and, more important, a fix for the ”core dump” bug when set up as a secondary name server (well, now everything is linked static); some cleanups were done to the Makefiles; this is just an update release, it contains only executables and this README!

5 To do

• either replace makefiles with REXX scripts or craft’em right

• include a working sample configuration

• find and fix bugs in the port :)

• make named display some debug info in its window

• provide some means to send signals to named

• figure out a way to have dynamic linking and a working fork()

6 How to contact the author

I’m very interested in your comments and flames.

Peter Meerwald pmeerw@cosy.sbg.ac.at Schiestandstr. 3 Peter.Meerwald@2-315-11-0.st.co.at A-5061 Elsbethen/AUSTRIA 2:315/11@fidonet.org

7 Disclaimer

Beware: I am no DNS specialist. There may be serious bugs in the OS/2 port of bind 4.9.3 – use at your own risk! Insert your favourite disclaimer text below…

6a unique number identifying a process

7again on hobbes or http://www.apache.org/